# On the Alexander polynomial of special alternating links

### v2

@@ 6.1. Crowell’s model
The following combinatorial model for the Alexander polynomial of alternating links is due to Crowell [3]. Recall that a link is alternating if it has an alternating diagram. A diagram is alternating if its undercrossings and overcrossings alternate as we trace it.
Let $\mathcal{G}(L)$ be the planar graph obtained by flattening the crossings of an alternating diagram of $L$: the crossings of $L$ are the vertices of $\mathcal{G}(L)$, and the arcs between the crossings are the edges of $\mathcal{G}(L)$. Note that $\mathcal{G}(L)$ is a planar $4$-regular $2$-face colorable graph. Next, we assign directions and weights to the edges of $\mathcal{G}(L)$ in the following way:
@@
On the left, we see the orientation of the link $L$ in an overcrossing, and on the right, we see how the edges of $\mathcal{G}(L)$ are directed and weighted. Note that the edge orientations assigned to $\overrightarrow{\mathcal{G}(L)}$ are not the same as the orientation of the corresponding components of the link. Denote by $\overrightarrow{\mathcal{G}(L)}$ the oriented weighted graph obtained from $\mathcal{G}(L)$ in this fashion. Let $var(e)$ denote the weight $-t$ or $1$ assigned to the edge $e\in E(\overrightarrow{\mathcal{G}(L)})$. See Figure 8 for a full example.
@@
Given an alternating diagram of the link $L$, fix an arbitrary vertex $v\in V(\overrightarrow{\mathcal{G}(L)})$. The Alexander polynomial of $L$ is:
$\Delta_{L}(t)\sim\sum_{\begin{subarray}{c}A\in\mathcal{A}_{v}(\overrightarrow{\mathcal{G}(L)})\end{subarray}}\prod_{e\in E(A)}var(e).$
Remark 6.2.
[…]
@@ 6.2. Definition of our weight-preserving bijection
Fix $v\in\operatorname{V}(\overrightarrow{\mathcal{G}(L)})$ and $f_{v}\in\operatorname{E}(\overrightarrow{\mathcal{G}(L)})$ with final vertex $v$. Let $\mathcal{C}(\overrightarrow{\mathcal{G}(L)})$ denote the set of Eulerian tours of $\overrightarrow{\mathcal{G}(L)}$, and let $\mathcal{T}(D)$ denote the set of spanning trees of $D$. Recall that without its edge orientation, $\overrightarrow{\mathcal{G}(L)}$ is the medial graph of $G$ (see Section 1.2). Since $D$ is the planar dual of $G$, $\overrightarrow{\mathcal{G}(L)}$ can equivalently be constructed as the medial graph of $D$. In this section, we will write $v_{e}$ to mean the vertex of $\overrightarrow{\mathcal{G}(L)}$ corresponding to an edge $e\in E(D)$, as in this medial graph construction.
@@
Let $\Phi:\mathcal{T}(D)\rightarrow\mathcal{C}(\overrightarrow{\mathcal{G}(L)})$ be the map which, given $T\in\mathcal{T}(D)$, produces $\Phi(T)$ by the following procedure. At each vertex $v_{e}$ of $\overrightarrow{\mathcal{G}(L)}$, we choose a pairing of incident edges as shown below, depending on whether or not $e$ is in $E(T)$.
On the left, we show an edge $e\in E(D)$ meeting the vertex $v_{e}$ in $\overrightarrow{\mathcal{G}(L)}$. The center diagram shows the choice of edge pairing in $\overrightarrow{\mathcal{G}(L)}$ when $e\in E(T)$, and the right diagram shows the choice of edge pairing in $\overrightarrow{\mathcal{G}(L)}$ when $e\notin E(T)$.
Figure 9 depicts an example of a spanning tree and its image under $\Phi$ interpreted as an Eulerian tour of $\overrightarrow{\mathcal{G}(L)}$.
@@
An eloquent exposition of the map from Definition 6.4 can be found in Stanley’s account of the BEST Theorem [19]. Applying [19, Theorem 10.2] to $\overrightarrow{\mathcal{G}(L)}^{T}$ with the additional condition that every vertex of $\overrightarrow{\mathcal{G}(L)}$ has outdegree $2$, we readily obtain that $\Psi_{v,f_{v}}$ is a bijection. Recall that $\overrightarrow{\mathcal{G}(L)}^{T}$ denotes the transpose of $\overrightarrow{\mathcal{G}(L)}$, i.e. the digraph obtained by reversing the direction of each edge in $\overrightarrow{\mathcal{G}(L)}$. Refer back to Section 2 for more details on the BEST theorem.
The map $F=\Psi_{v,f_{v}}\circ\Phi$ is a bijection between spanning trees of $D$ and arborescences of $\overrightarrow{\mathcal{G}(L)}$. In fact, Theorem 6.5 says a lot more: $F$ is weight preserving (up to multiplication by a fixed monomial) when the spanning trees of $D$ and arborescences of $\overrightarrow{\mathcal{G}(L)}$ are weighted as in the Kauffman and Crowell models, respectively. For fixed $r\in\operatorname{V}(D)$ and for $T\in\mathcal{T}(D)$, we will write $\text{wt}_{K}(T)=t^{|\operatorname{E}(T)\setminus\operatorname{E}_{\text{away}}(T)|}$ for the Kauffman weight, as explained in Remark 4.4, where $\operatorname{E}_{\text{away}}(T)$ is the subset of edges of $T$ such that reversing the orientation of every edge in $\operatorname{E}_{\text{away}}(T)$ would make $T$ an oriented spanning tree rooted at $r$. Specifically, we leverage Lemma 4.2 and omit the contribution of the spanning tree of $G$ to each monomial. For fixed $v\in\operatorname{V}(\overrightarrow{\mathcal{G}(L)})$ and $A\in\mathcal{A}_{v}(\overrightarrow{\mathcal{G}(L)})$, we write $\text{wt}_{C}(A)=\prod_{e\in\operatorname{E}(A)}var(e)$ for the Crowell weight of an arborescence, as in Remark 6.2.
To make sense of the notation $m_{2}$ used in the statement of Theorem 6.5, we set the following convention. Refer to the clockwise oriented regions of $\overrightarrow{\mathcal{G}(L)}$ by $R(\overrightarrow{\mathcal{G}(L)})$. It was shown in [5, Lemma 3.1] that for each of these regions, the edges are either labeled exclusively by $-t$ or exclusively by $1$ in Crowell’s state model. Moreover, for any pair of adjacent vertices in $V(G)$, one of the corresponding regions must be labeled with $-t$ and the other with $1$. We can thus sort the elements of $R(\overrightarrow{\mathcal{G}(L)})$ into color classes $\{R_{1},\dots,R_{m_{1}}\}\sqcup\{R_{\overline{1}},\dots,R_{\overline{m_{2}}}\}$ where $\{R_{1},\dots,R_{m_{1}}\}$ are the regions whose boundary edges are labeled by $-t$ in Crowell’s state model. We refer to $\{R_{1},\dots,R_{m_{1}}\}$ and $\{R_{\overline{1}},\dots,R_{\overline{m_{2}}}\}$ as gold and blue regions respectively.
@@
Let $r\in\operatorname{V}(D)$ denote the vertex corresponding to the exterior region of $L$. Choose $v\in\operatorname{V}(\overrightarrow{\mathcal{G}(L)})$ such that $v=v_{e}$ for some $e\in E(D)$ with $\operatorname{fin}(e)=r$, and let $f_{v}\in\operatorname{E}(\overrightarrow{\mathcal{G}(L)})$ be the unique edge in the boundary of the exterior region such that $\operatorname{fin}(f_{v})=v$.
@@
Figure 10 depicts an instance of the above bijection.
@@
To prove Theorem 6.5, we introduce two lemmas and accompanying notation.
@@
Fix $T\in\mathcal{T}(D)$. For $e\in E(T)$ such that $v_{e}\neq v$, the root of $F(T)$, let $f_{e}\in\operatorname{E}(F(T))$ denote the unique edge of the arborescence $F(T)$ with final vertex $v_{e}$ (see Figure 11).
@@
Fix a spanning tree $T$ of $D$, and let $r\in\operatorname{V}(D)$ denote the vertex corresponding to the exterior region of $L$. Fix $v\in\operatorname{V}(\overrightarrow{\mathcal{G}(L)})$ such that $v=v_{e}$ for some $e\in E(D)$ with $\operatorname{fin}(e)=r$. Let $f_{v}\in\operatorname{E}(\overrightarrow{\mathcal{G}(L)})$ be the unique edge in the boundary of the exterior region such that $\operatorname{fin}(f_{v})=v$. Then the edges in $\{f_{e}\ |\ e\in\operatorname{E}(T)\ \text{points towards}\ r\text{ in $T$}\}$ all border gold regions of $\overrightarrow{\mathcal{G}(L)}$, and the edges in $\{f_{e}\ |\ e\in\operatorname{E}(T)\ \text{points away from}\ r\text{ in $T$}\}$ all border blue regions of $\overrightarrow{\mathcal{G}(L)}$.
@@
For ease of reference, we say an edge of $\overrightarrow{\mathcal{G}(L)}$ is gold (resp. blue) if it borders a gold (resp. blue) region. Fix a spanning tree $T$ of $D$. We outline a procedure to determine the color of the first incoming edge to $v_{e}$ in $\Phi(T)$ for each $e\in\operatorname{E}(T)$. Note that the first incoming edge to $v_{e}$ in $\Phi(T)$ for each $e\in\operatorname{E}(T)$ is indeed $f_{e}$ according to Definitions 6.4 and 6.6.
Suppose first that $e\in\operatorname{E}(T)$ has $r$ as an endpoint. Let $P$ denote the portion of the oriented boundary of the exterior region of $\overrightarrow{\mathcal{G}(L)}$ from $v$ to $v_{e}$. We claim that the last edge of $P$ is the first edge incoming to $v_{e}$ along $\Phi(T)$ (which is $f_{e}$). Note that $\Phi(T)$, starting at $f_{v}$, can only leave the boundary of the exterior region when it meets a vertex of the form $v_{e^{\prime}}$ for some $e^{\prime}\in\operatorname{E}(T)$. Note also that upon leaving the boundary of the exterior region from a vertex of the form $v_{e^{\prime}}$ for some $e^{\prime}\in\operatorname{E}(T)$, the Eulerian cycle $\Phi(T)$ will return to $v_{e^{\prime}}$ before meeting any other vertex $v_{e^{\prime\prime}}$, $e^{\prime\prime}\in E(T)$, in the boundary of the exterior region. If the opposite were true, $T$ would contain a cycle, contradicting the fact that $T$ is a tree.
Accordingly, the first edge incident to $v_{e}$ along $\Phi(T)$ is blue if and only if the last edge of $P$ is blue. This happens if and only if $e$ is directed away from $r$ since the edges incident to $r$ alternate between incoming and outgoing (see the left and center images in Figure 12).
Figure 12. The procedure described in the proof of Lemma 6.7 to determine the first edge incident to $v_{e_{i}}$ for each of the edges of the spanning tree from Figures 10 and 11.
Next, we consider the case that $e\in\operatorname{E}(T)$ is not incident to $r$. Fix the shortest (undirected) path in $T$ from $r$ to a vertex of $e$, and let $e^{\prime\prime\prime}\neq e$ denote the final edge of this path. The edges $e$ and $e^{\prime\prime\prime}$ meet at a single vertex $w$. Let $P^{\prime}$ denote the oriented path from $v_{e^{\prime\prime\prime}}$ to $v_{e}$ in the boundary of the region of $\overrightarrow{\mathcal{G}(L)}$ corresponding to $w$ (in the rightmost image of Figure 12, $e$ is $e_{3}$, $e^{\prime\prime\prime}$ is $e_{2}$, and $w$ is the vertex where they meet). As before, the last edge of $P^{\prime}$ is the first edge incident to $v_{e}$ along $\Phi(T)$. If the length of the path connecting $v_{e^{\prime\prime\prime}}$ and $v_{e}$ is odd, then the colors of the first edges incident to $v_{e^{\prime\prime\prime}}$ and $v_{e}$ are identical, and $e^{\prime\prime\prime}$ and $e$ either both point towards $v$ or both point away from $v$. Similarly, if the length of the path connecting $v_{e^{\prime\prime\prime}}$ and $v_{e}$ is even, then the colors of the first edges incident to $v_{e^{\prime\prime\prime}}$ and $v_{e}$ are different, and exactly one of these edges points towards $v$. We, thus, achieve the desired result. ∎
@@
Fix a spanning tree $T$ of $D$, and let $r\in\operatorname{V}(D)$ denote the exterior region of $L$. Fix $v\in\operatorname{V}(\overrightarrow{\mathcal{G}(L)})$ such that $v=v_{e}$ for some $e\in E(D)$ with $\operatorname{fin}(e)=r$. Let $f_{v}\in\operatorname{E}(\overrightarrow{\mathcal{G}(L)})$ be the unique edge in the boundary of the exterior region such that $\operatorname{fin}(f_{v})=v$. Let $f\in\operatorname{E}(\overrightarrow{\mathcal{G}(L)})$ be any edge bordering a blue region $C_{f}$ of $\overrightarrow{\mathcal{G}(L)}$. Then $f\in\operatorname{E}(F(T))$ if and only if
  • a.
    $f$ is not the last edge in the boundary of $C_{f}$ to appear in the Eulerian tour $\Phi(T)$, and
  • b.
    the final vertex $\operatorname{fin}(f)$ in $\overrightarrow{\mathcal{G}(L)}$ does not equal $v_{e}$ for some edge $e\in\operatorname{E}(T)$ directed towards the root $r$.
@@
First, suppose that $f$ is the last edge in the boundary of $C_{f}$ to appear in $\Phi(T)$. Then $f$ is first edge pointing into $\operatorname{fin}(f)$ in $\Phi(T)$ only when $\operatorname{fin}(f)=r$, so $f\notin\operatorname{E}(F(T))$.
If $\operatorname{fin}(f)=v_{e}$ where $e\in\operatorname{V}(T)$ is directed towards the root $r$, then by Lemma 6.7, $f_{e}$ must lie in the boundary of a gold region of $\overrightarrow{\mathcal{G}(L)}$. In particular, $f\neq f_{e}$, and so $f$ is not an edge of $F(T)$.
It now remains to show that all other edges $f$ which border a blue region of $\overrightarrow{\mathcal{G}(L)}$ must be in $F(T)$. Suppose $f$ satisfies conditions (a) and (b). One of the following cases must hold.
  1. Case 1: The final vertex $\operatorname{fin}(f)=v_{e}$ for some edge $e\in\operatorname{E}(T)$ directed away from the root $r$. (Note that this implies $f\neq f_{v}$.)
    In this case, Lemma 6.7 implies that $f$ is an edge of $F(T)$.
@@
  2. Case 2: The final vertex $\operatorname{fin}(f)=v_{e}$ for some edge $e\in\operatorname{E}(D)$ which is not in $T$, and the unique edge $f^{\prime}$ in the boundary of $C_{f}$ with $\operatorname{fin}(f)=\operatorname{init}(f^{\prime})$ appears after $f$ in the Eulerian tour $\Phi(T)$.

### v3

@@ 6.1. Crowell’s model
The following combinatorial model for the Alexander polynomial of alternating links is due to Crowell [4]. Recall that a link is alternating if it has an alternating diagram. A link diagram is alternating if its undercrossings and overcrossings alternate as we trace along each component.
Let $\mathcal{G}(L)$ be the planar graph obtained by flattening the crossings of an alternating diagram of $L$: the crossings of $L$ are the vertices of $\mathcal{G}(L)$, and the arcs between the crossings are the edges of $\mathcal{G}(L)$. Note that $\mathcal{G}(L)$ is a planar $4$-regular graph. Next, we assign directions and weights to the edges of $\mathcal{G}(L)$ in the following way:
@@
On the left, we see the orientation of the link $L$ at a crossing, and on the right, we see how the edges of $\mathcal{G}(L)$ incident to the corresponding vertex are directed and weighted. Denote by $\overrightarrow{\mathcal{G}(L)}$ the oriented weighted graph obtained from $\mathcal{G}(L)$ in this fashion. Note that the edge orientations assigned to $\overrightarrow{\mathcal{G}(L)}$ are not the same as the orientation of the corresponding components of the link. Let $\operatorname{var}(e)$ denote the weight $-t$ or $1$ assigned to the edge $e\in\operatorname{E}(\overrightarrow{\mathcal{G}(L)})$. See Figure 8 for a full example.
@@
Notation 6.1.
For every $A\in\mathcal{A}_{v}(\overrightarrow{\mathcal{G}(L)})$, we define the Crowell weight to be
$\operatorname{wt}_{C}(A)=\prod_{e\in E(A)}\operatorname{var}(e).$
@@
Given an alternating diagram of the link $L$, fix an arbitrary vertex $v\in\operatorname{V}\overrightarrow{\mathcal{G}(L)})$. The Alexander polynomial of $L$ is:
$\Delta_{L}(t)\sim\sum_{\begin{subarray}{c}A\in\mathcal{A}_{v}(\overrightarrow{\mathcal{G}(L)})\end{subarray}}\operatorname{wt}_{C}(A).$
@@ 6.2. Definition of our weight-preserving bijection
The goal of this section is to define a bijection between arborescences of $\overrightarrow{\mathcal{G}(L)}$ and spanning trees of $D$ that is weight-preserving up to multiplication by a fixed monomial. For fixed $v\in\operatorname{V}(\overrightarrow{\mathcal{G}(L)})$ and $A\in\mathcal{A}_{v}(\overrightarrow{\mathcal{G}(L)})$, recall that the Crowell weight of an arborescence $A$ is the monomial $\operatorname{wt}_{C}(A)=\prod_{e\in\operatorname{E}(A)}\operatorname{var}(e)$. We define the Kauffman weight of a spanning tree of $D$ rooted at $r\in\operatorname{V}(D)$ as follows.
Notation 6.3.
For fixed $r\in\operatorname{V}(D)$ and for $T\in\mathcal{T}(D)$, we will write $\operatorname{wt}_{K}(T)=t^{|\operatorname{E}(T)\setminus\operatorname{E}_{\text{away}}(T)|}$ for the Kauffman weight, as explained in Remark 4.4, where $\operatorname{E}_{\text{away}}(T)$ is the subset of edges of $T$ such that reversing the orientation of every edge in $\operatorname{E}_{\text{away}}(T)$ would make $T$ an oriented spanning tree rooted at $r$.
Fix $v\in\operatorname{V}(\overrightarrow{\mathcal{G}(L)})$ in the boundary of the exterior region of $\overrightarrow{\mathcal{G}(L)}$, and let $f_{v}\in\operatorname{E}(\overrightarrow{\mathcal{G}(L)})$ be the unique edge in the boundary of the exterior region with final vertex $v$. Let $\mathcal{C}(\overrightarrow{\mathcal{G}(L)})$ denote the set of Eulerian tours of $\overrightarrow{\mathcal{G}(L)}$ considered as paths beginning with $f_{v}$. Let $\mathcal{T}(D)$ denote the set of spanning trees of $D$. Recall that $\mathcal{G}(L)$ is the medial graph of $G$ (see Section 1.2). Since $D$ (considered without its orientation) is the planar dual of $G$, $\overrightarrow{\mathcal{G}(L)}$ can equivalently be constructed by assigning edge weights and orientations to the medial graph of $D$. In this section, we will write $v_{e}$ to mean the vertex of $\overrightarrow{\mathcal{G}(L)}$ corresponding to an edge $e\in\operatorname{E}(D)$, as in this medial graph construction.
@@
Let $\Phi:\mathcal{T}(D)\rightarrow\mathcal{C}(\overrightarrow{\mathcal{G}(L)})$ be the map which, given $T\in\mathcal{T}(D)$, produces $\Phi(T)$ by the following procedure. At each vertex $v_{e}$ of $\overrightarrow{\mathcal{G}(L)}$, we choose a pairing of incident edges as shown below, depending on whether or not $e$ is in $\operatorname{E}(T)$.
On the left, we show an edge $e\in\operatorname{E}(D)$ meeting the vertex $v_{e}$ in $\overrightarrow{\mathcal{G}(L)}$. The center diagram shows the choice of edge pairing in $\overrightarrow{\mathcal{G}(L)}$ when $e\in\operatorname{E}(T)$, and the right diagram shows the choice of edge pairing in $\overrightarrow{\mathcal{G}(L)}$ when $e\notin\operatorname{E}(T)$.
Figure 9 depicts an example of a spanning tree and its image under $\Phi$.
@@
An eloquent exposition of the map from Definition 6.5 can be found in Stanley’s account of the BEST Theorem [21]. Recall that $\overrightarrow{\mathcal{G}(L)}^{T}$ denotes the transpose of $\overrightarrow{\mathcal{G}(L)}$, i.e. the digraph obtained by reversing the direction of each edge in $\overrightarrow{\mathcal{G}(L)}$. Applying [21, Theorem 10.2] to $\overrightarrow{\mathcal{G}(L)}^{T}$ with the additional condition that every vertex of $\overrightarrow{\mathcal{G}(L)}$ has outdegree $2$, we readily obtain that $\Psi_{v,f_{v}}$ is a bijection. Refer back to Section 2 for more details on the BEST theorem.
The map $F=\Psi_{v,f_{v}}\circ\Phi$ is a bijection between spanning trees of $D$ and arborescences of $\overrightarrow{\mathcal{G}(L)}$. In fact, Theorem 6.7 says a lot more: $F$ is weight-preserving (up to multiplication by a fixed monomial which depends only on graph-theoretic properties of $D$) when the spanning trees of $D$ and arborescences of $\overrightarrow{\mathcal{G}(L)}$ are weighted as in Notation 6.3 and Notation 6.1, respectively.
We set the following convention to make sense of the notation $m_{2}$ used in the statement of Theorem 6.7. Refer to the clockwise oriented regions of $\overrightarrow{\mathcal{G}(L)}$ by $R(\overrightarrow{\mathcal{G}(L)})$. It was shown in [6, Lemma 3.1] that for each of these regions, the edges are either labeled exclusively by $-t$ or exclusively by $1$ in Crowell’s model. Moreover, for any pair of adjacent vertices in $\operatorname{V}(G)$, one of the corresponding regions must be labeled with $-t$ and the other with $1$. We can thus sort the elements of $R(\overrightarrow{\mathcal{G}(L)})$ into color classes $\{R_{1},\dots,R_{m_{1}}\}\sqcup\{R_{\overline{1}},\dots,R_{\overline{m_{2}}}\}$ where $\{R_{1},\dots,R_{m_{1}}\}$ are the regions whose boundary edges are labeled by $-t$ in Crowell’s model. We refer to $\{R_{1},\dots,R_{m_{1}}\}$ and $\{R_{\overline{1}},\dots,R_{\overline{m_{2}}}\}$ as gold and blue regions respectively.
Remark 6.6.
Let $G$ be a planar bipartite graph. Comparing the definitions of gold and blue with Figure 5, a region in $R(\overrightarrow{\mathcal{G}(L_{G})})$ is gold if and only if the corresponding region of $L_{G}$ has clockwise oriented boundary.
@@
Let $r\in\operatorname{V}(D)$ denote the vertex corresponding to the exterior region of $L$. Fix $e_{0}\in\operatorname{E}(D)$ with $\operatorname{fin}(e_{0})=r$. Let $v=v_{e_{0}}\in\operatorname{V}(\overrightarrow{\mathcal{G}(L)})$, and let $f_{v}\in\operatorname{E}(\overrightarrow{\mathcal{G}(L)})$ be the unique edge in the boundary of the exterior region such that $\operatorname{fin}(f_{v})=v$.
@@
Figure 10 depicts an instance of the bijection $F$.
@@
To prove Theorem 6.7, we introduce four lemmas and accompanying notation.
@@
Fix $T\in\mathcal{T}(D)$. For $e\in\operatorname{E}(T)$ such that $v_{e}\neq v$, the root of $F(T)$, let $f_{e}\in\operatorname{E}(F(T))$ denote the unique edge of the arborescence $F(T)$ with final vertex $v_{e}$ (see Figure 11). In particular, $f_{e}$ is the first edge incoming to $v_{e}$ in $\Phi(T)$.
@@
Fix $T\in\mathcal{T}(D)$, let $w\in\operatorname{V}(D)$, and let $e_{1}\in\operatorname{E}(T)$ be incident to $w$. Let $f=f_{e_{1}}$ in $\Phi(T)$ and let $g$ be the other edge of $\overrightarrow{\mathcal{G}(L)}$ with final vertex $v_{e_{1}}$. If $f$ is in the boundary of the region of $\overrightarrow{\mathcal{G}(L)}$ corresponding to $w$, then for any $e_{2}\in\operatorname{E}(T)$ with vertex $w$ such that $e_{2}\neq e_{1}$, the portion of $\Phi(T)$ between the edges $f$ and $g$ does not meet $v_{e_{2}}$.
@@
Observe that if $e^{\prime},e^{\prime\prime}\in\operatorname{E}(T)$ are such that $\Phi(T)$ reaches $v_{e^{\prime}}$ before $v_{e^{\prime\prime}}$ and no other vertex in the path between $v_{e^{\prime}}$ and $v_{e^{\prime\prime}}$ in $\Phi(T)$ has the form $v_{e^{\prime\prime\prime}}$ for any $e^{\prime\prime\prime}\in\operatorname{E}(T)$, then $e^{\prime}$ and $e^{\prime\prime}$ share an endpoint in $D$. In particular, $e^{\prime}$ and $e^{\prime\prime}$ are either adjacent or equal.
Let $v_{e_{1}},v_{e^{\prime}_{1}}\ldots,v_{e^{\prime}_{m}},v_{e_{1}}$ be the set of all vertices of the form $v_{e^{\prime}}$ for $e^{\prime}\in\operatorname{E}(T)$ which are reached by $\Phi(T)$ in between the edges $f$ and $g$. Suppose (in search of a contradiction) that $v_{e^{\prime}_{k}}$ has $w$ as an endpoint for some $1\leq k\leq m$. By the observation above, we can see that $e_{1},e^{\prime}_{1},\ldots,e^{\prime}_{k}$ is a path from $w$ back to $w$ within $T$, and since $e_{1}\neq e^{\prime}_{k}$, this path must contain a cycle. This contradicts the fact that $T$ is a tree. We thus conclude that the lemma holds. ∎
Given a spanning tree $T$ of $D$, understanding the weight of $F(T)$ requires one to know precisely how many edges of $F(T)$ border blue regions of $\overrightarrow{\mathcal{G}}(L)$. Lemma 6.10 concerns the edges of the form $f_{e}$ for some $e\in\operatorname{E}(T)$, and Lemma 6.11 concerns the remaining edges.
@@
Fix a spanning tree $T$ of $D$, and let $r\in\operatorname{V}(D)$ denote the vertex corresponding to the exterior region of $\overrightarrow{\mathcal{G}(L)}$. Fix $e_{0}\in\operatorname{E}(D)$ with $\operatorname{fin}(e_{0})=r$. Let $v=v_{e_{0}}\in\operatorname{V}(\overrightarrow{\mathcal{G}(L)})$, and let $f_{v}\in\operatorname{E}(\overrightarrow{\mathcal{G}(L)})$ be the unique edge in the boundary of the exterior region such that $\operatorname{fin}(f_{v})=v$.
For $e\in\operatorname{E}(T)$, $f_{e}$ is blue if and only if $e\in\operatorname{E}_{\text{away}}(T)$.
@@
The lemma statement is equivalent to saying the edges in $\{f_{e}\ |\ e\in\operatorname{E}(T)\ \text{points towards}\ r\text{ in $T$}\}$ all border gold regions of $\overrightarrow{\mathcal{G}(L)}$, and the edges in $\{f_{e}\ |\ e\in\operatorname{E}(T)\ \text{points away from}\ r\text{ in $T$}\}$ all border blue regions of $\overrightarrow{\mathcal{G}(L)}$. For ease of reference, we say an edge of $\overrightarrow{\mathcal{G}(L)}$ is gold (resp. blue) if it borders a gold (resp. blue) region.
Fix a spanning tree $T$ of $D$. Recall that for every $e\in\operatorname{E}(T)$, Definitions 6.5 and 6.8 state that $f_{e}$ is the first incoming edge to $v_{e}$ in $\Phi(T)$, where we treat $\Phi(T)$ as beginning with the edge $f_{v}$. We outline the following procedure to determine the color of $f_{e}$ for each $e\in\operatorname{E}(T)$.
  • Case 1: The edge $e$ of $T$ has $r$ as an endpoint.
    By definition, this is equivalent to the case that $v_{e}$ is in the boundary of the exterior region of $\overrightarrow{\mathcal{G}(L)}$. Let $P$ denote the portion of the oriented boundary of the exterior region of $\overrightarrow{\mathcal{G}(L)}$ from $v$ to $v_{e}$ (see the left and center images of Figure 12). We claim that the last edge of $P$ is the first edge incoming to $v_{e}$ along $\Phi(T)$ (which is $f_{e}$). Note that $\Phi(T)$, starting at $f_{v}$, can only leave the boundary of the exterior region when it meets a vertex of the form $v_{e^{\prime}}$ for some $e^{\prime}\in\operatorname{E}(T)$. Note also that upon leaving the boundary of the exterior region from a vertex of the form $v_{e^{\prime}}$ for some $e^{\prime}\in\operatorname{E}(T)$, Lemma 6.9 states that the Eulerian cycle $\Phi(T)$ will return to $v_{e^{\prime}}$ before meeting any other vertex $v_{e^{\prime\prime}}$, $e^{\prime\prime}\in\operatorname{E}(T)$, in the boundary of the exterior region.
    Accordingly, the first edge incident to $v_{e}$ along $\Phi(T)$ is blue if and only if the last edge of $P$ is blue. This happens if and only if $e$ is directed away from $r$ since the edges incident to $r$ alternate between incoming and outgoing. The left and center images in Figure 12 show examples of this case.
    Figure 12. The procedure described in the proof of Lemma 6.10 to determine the first edge incident to $v_{e_{i}}$ for each of the edges of the spanning tree from Figures 10 and 11. In the left and center images, the path $P$ is indicated in blue. In the right image, the blue line indicates the path $P^{\prime}$ from $v_{e_{2}}$ to $v_{e_{3}}$ appended to the path $P$ from $v$ to $v_{e_{2}}$.
  • Case 2: The edge $e\in\operatorname{E}(T)$ is not incident to $r$.
    Fix the shortest (undirected) path in $T$ from $r$ to a vertex of $e$, and let $e^{\prime}\neq e$ denote the final edge of this path. The edges $e$ and $e^{\prime}$ meet at a single vertex $w$. Let $P^{\prime}$ denote the oriented path from $v_{e^{\prime}}$ to $v_{e}$ in the boundary of the region of $\overrightarrow{\mathcal{G}(L)}$ corresponding to $w$ (in the rightmost image of Figure 12, $e$ is $e_{3}$, $e^{\prime}$ is $e_{2}$, and $w$ is the vertex where they meet). As in Case 1, Lemma 6.9 implies that the last edge of $P^{\prime}$ is the first edge incident to $v_{e}$ along $\Phi(T)$. If the length of the path connecting $v_{e^{\prime}}$ and $v_{e}$ is odd, then since the colors of the edges surrounding any region of $\overrightarrow{\mathcal{G}(L)}$ bounded by a counterclockwise cycle alternate and $D$ is an alternating dimap, the colors of the first edges incident to $v_{e^{\prime}}$ and $v_{e}$ are identical, and $e^{\prime}$ and $e$ either both point towards $v$ or both point away from $v$. Similarly, if the length of the path connecting $v_{e^{\prime}}$ and $v_{e}$ is even, then the colors of the first edges incident to $v_{e^{\prime}}$ and $v_{e}$ are different, and exactly one of these edges points towards $v$. We thus achieve the desired result.
Lemma 6.11.
Fix a spanning tree $T$ of $D$, and let $r\in\operatorname{V}(D)$ denote the vertex corresponding to the exterior region of $L$. Fix $e_{0}\in\operatorname{E}(D)$ with $\operatorname{fin}(e_{0})=r$. Let $v=v_{e_{0}}\in\operatorname{V}(\overrightarrow{\mathcal{G}(L)})$, and let $f_{v}\in\operatorname{E}(\overrightarrow{\mathcal{G}(L)})$ be the unique edge in the boundary of the exterior region such that $\operatorname{fin}(f_{v})=v$.
Let $f$ be a blue edge of $\overrightarrow{\mathcal{G}(L)}$, and let $R$ denote the blue planar region it bounds. Suppose $\operatorname{fin}(f)\neq v_{e}$ for any $e\in\operatorname{E}(T)$. Then $f$ is an edge of $F(T)$ if and only if $f$ is not the last edge in the boundary of $R$ to appear in $\Phi(T)$.
Proof.
@@
First, suppose that $f$ is the last edge in the boundary of $R$ to appear in $\Phi(T)$. Then $f$ is the first edge pointing into $\operatorname{fin}(f)$ in $\Phi(T)$ only when $\operatorname{fin}(f)=v$, so $f\notin\operatorname{E}(F(T))$.
Next, suppose $f$ is not the last edge in the boundary of $R$ reached by $\Phi(T)$. Let $f^{\prime}$ be the unique edge such that $\operatorname{init}(f^{\prime})=\operatorname{fin}(f)$. There are two possibilities.
  1. (1)
    Either the edge $f^{\prime}$ appears before $f$ in $\Phi(T)$, or
  2. (2)
    the edge $f^{\prime}$ appears after $f$ in $\Phi(T)$.
We argue that in the first case, $f\in\operatorname{E}(F(T))$ and that the second case is impossible.
@@
  1. Case 1: The final vertex $\operatorname{fin}(f)=v_{e}$ for some edge $e\in\operatorname{E}(D)$ which is not in $T$, and the unique edge $f^{\prime}$ in the boundary of $R$ with $\operatorname{fin}(f)=\operatorname{init}(f^{\prime})$ appears after $f$ in the Eulerian tour $\Phi(T)$.